OpenStudy (anonymous):

A rectangle has a width of 5 inches. The length is 30% more than the width. Find the perimeter of the rectangle.

OpenStudy (anonymous):

23

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

L = 30% more then the width L = 5 + .30(5) = 5 + 1.5 = 6.5 P = 2L + 2W P = 2(6.5) + 2(5) P = 13 + 10 P = 23

OpenStudy (texaschic101):

just thought you might want to know how to get the answer

OpenStudy (anonymous):

thank you it makes a lot more sense now